OpenStudy (anonymous):

Divide and check. 49x^7-56x^5+42x^3/7x^3

OpenStudy (anonymous):

49x^7-56x^5+6

OpenStudy (anonymous):

factor the top and cacel the bottom

OpenStudy (anonymous):

doing that should leave you with 7x^4-8x^2+6

OpenStudy (anonymous):

7x^\[7x^{3}(7x^4-8x^2+6) \over7x^3\]

OpenStudy (anonymous):

7x^3 cancel

OpenStudy (anonymous):

leaving you with 7x^4-8x^2+6

OpenStudy (agreene):

\[\frac{49x^7-56x^5+42x^3}{7x^3}\] \[=\frac{49x^7}{7x^3}-\frac{56x^5}{7x^3}+\frac{42x^3}{7x^2}\] \[=7x^4-8x^2+6\]

OpenStudy (anonymous):

i am not sure how you guys can type math expressions

OpenStudy (anonymous):

the equation editor in the bottom

OpenStudy (agreene):

\[\huge skill\]

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

well THX all

OpenStudy (agreene):

I dont use the eqn editor, I just type 'em out \[\large \frac{\rho}{\Delta}\int F'(x)dx\] and such